# Designing a skilled soccer team for RoboCup: exploring Skill-Set-Primitives through reinforcement learning

### v1

@@ 6. Conversion
Reset
The robot is reset to a fixed initial pose with bent knees, which is then added as a continuous bias (see Table 5, 5a)
Termination
Robot has fallen ($z<0.27$) or timeout (4 s)
Episode
There is no specific configuration during the episode
Reward
$r=\frac{d\text{x}}{dt}\,$ for $\Delta t=0.06$ s. There is no penalty for deviations in $y$. The reward is updated every vision cycle.
@@
Reset
Sprint Forward is executed for 1.4 s (this parameter was found to be the minimal period for which the resulting behavior, Sprint & Rotate, achieves maximum speed)
Termination
Robot has fallen ($z<0.34$) or timeout (7 s)
Episode
A new target direction is randomly generated every 0.28 s (cycle caption period, see Fig. 3). The target’s range is [-10,10] deg, relative to the current orientation of the robot’s torso.
Reward
@@
Reset
The ball is placed in a random but reachable position. Sprint Forward is executed for 1.4 s. Then, Sprint & Rotate is executed until the ball is within 1 meter.
Termination
Robot has fallen ($z<0.33$) or timeout (0.4 s)
Episode
There is no specific configuration during the episode
Reward
@@
Table 9: Sprint-Kick hyperparameters: Sprint Forward (SF), Sprint & Rotate (SR), Kick (K)
@@ A.3 Locomotion set
step
1
Step counter (max: 120)a [3]
0
1
z
1
Head heightb
0
1
dz
1
$\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s
0
1
roll
1
Torso roll angleb
0
-1
pitch
1
Torso pitch angleb
0
1
gyro
3
Gyroscope (x-forward, y-left, z-up)
(0,1,2)
(-1,1,-1)
acc
3
Accelerometer (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
feet
12
(l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
(6,7,8,9,10,11, 0,1,2,3,4,5)
(1,-$,1$,-1,1,1-1,1)
ankles
6
(l)eft and (r)ight ankle position relative to the robot’s hip: (lx,ly,lz,rx,ry,rz)
(3,4,5,0,1,2)
(1,-1,1,1,-1,1)
feetori
6
(l)eft and (r)ight feet orientation relative to the robot’s hip: (lx,ly,lz,rx,ry,rz)
(3,4,5,0,1,2)
(-$,-1$,1,-1)
shldrs
4
(l)eft and (r)ight shoulder joints positions
Sagittal reflectionc
Sagittal reflectionc
action
16
Action array generated in last time step
Sagittal reflectionc
Sagittal reflectionc
sbprog
1
Step Baseline cycle progress $\in[0,1]$
0
1
sbleg
2
One-hot encoding for the current Step Baseline swing leg (left leg, right leg)
(1,0)
(1,1)
targp
2
Filteredd 2D target position vector, relative to the robot’s torso (x-forward, y-left)
(0,1)
(1,-1)
targv
2
$\frac{d}{dt}\begin{bmatrix}\text{targp}_{x}&\text{targp}_{y}\end{bmatrix}$ for $\Delta t=0.02$ s
(0,1)
(1,-1)
targdir
1
Filteredd target direction, relative to the robot’s torso
0
-1
a A limit prevents the neural network from encountering higher values during testing than during training. The value 120 was optimized as a hyperparameter, as it influences the final behavior.
b Values extracted from 6D pose [4]
c The symmetry operation associated with this sagittal reflection depends on the joints’ order
d targp is limited to a magnitude of 0.5 m and a maximum variation of 0.014 m per time step (0.02 s), and targdir is limited to 45 deg and a maximum variation of 1.6 deg per time step
@@
1. Action
Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter
An exponential moving average filter is applied with a smoothing factor of 0.2
3. Global scale
The filtered action is multiplied by $\;\|\text{targp}\|_{2}\times 3.5+0.3\;$ if $\;\|\text{targp}\|_{2}<0.2\;$ to encourage the robot to minimize movement as it approaches the target position
4. Map
The individual values are mapped to the desired initial exploration range
5. Primitive
Add Step Baseline primitive
6. Constraints
Limit leg joints to prevent self-collisions
7. IK
Apply inverse kinematics to obtain the target joint positions
8. Conversion
First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.
@@ 8. Conversion
Reset
Execute Get Up behavior
Termination
Robot has fallen ($z<0.35$)
Episode
Two targets are created: position (TP) and direction (TD). A random velocity is periodically assigned (avg. interval of 1.6 s) to a moving TP within a work area of 196 $\text{m}^{2}$. Upon leaving the work area, the TP is relocated to its center. In 30% of episodes, TD coincides with TP, whereas in 70% of episodes, TD changes randomly at an avg. interval of 4 s.
Reward
@@
Parameter name Data size $\times$32b FP Description
step
1
Step countera [3]
z
1
Head heightb
dz
1
$\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s
roll
1
Torso roll angleb
pitch
1
Torso pitch angleb
gyro
3
Gyroscope (x-forward, y-left, z-up)
acc
3
Accelerometer (x-forward, y-left, z-up)
feet
12
(l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
joints
16
All joint positions excluding head, toes and 4 elbow joints
djoints
16
$\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{15}\end{bmatrix}$ for $\Delta t=0.02$ s
ballp
3
Ball position relative to the robot’s hip    (x-forward, y-left, z-up)
ballv
3
$\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s
balld
1
Euclidean distance from ball to robot’s hip
targdir
1
Target direction, relative to the robot’s torso
targd c
1
Target distance
a Step counter limits are not needed when testing episodes are never longer than training episodes
b Values extracted from 6D pose [4]
c The target distance is only included in Short Kick
@@
1. Action
Raw target speed for all joints, excluding head, toes and 4 elbow joints (16 values)
2. Map
The individual values are mapped to the desired initial exploration range
3. Bias
An initial bias is added to pull the kicking leg back for 0.04 s (for Short Kick) or 0.12 s (for Long Kick) to generate a back swing
@@ 3. Bias
Reset
The target direction is constant. The robot is initially placed near the ball with both a random position and orientation. The ball is given a random velocity with a magnitude inferior to 1 m/s. Then, the robot aligns itself with the ball and the target as it approaches. The alignment is stricter for the Long Kick. Additionally, for the Short Kick, the target distance is randomly set between 3 m and 9 m.
Termination
Timeout (0.20 s for Short Kick and 0.32 s for Long Kick)
Episode
There is no specific configuration during the episode
Reward
@@
step
1
Step counter (max: 96)a [3]
0
1
z
1
Head heightb
0
1
dz
1
$\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s
0
1
roll
1
Torso roll angleb
0
-1
pitch
1
Torso pitch angleb
0
1
gyro
3
Gyroscope (x-forward, y-left, z-up)
(0,1,2)
(-1,1,-1)
acc
3
Accelerometer (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
feet
12
(l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
(6,7,8,9,10,11, 0,1,2,3,4,5)
(1,-$,1$,-1,1,1-1,1)
joints
20
All joint positions excluding head and toes
Sagittal reflectionc
Sagittal reflectionc
djoints
20
$\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s
Sagittal reflectionc
Sagittal reflectionc
sbprog
1
Step Baseline cycle progress $\in[0,1]$
0
1
sbleg
2
One-hot encoding for the current Step Baseline swing leg (left leg, right leg)
(1,0)
(1,1)
sbsin
1
Step Baseline oscillation: sin(sbprog ×$\times\pi$)
0
1
ballp
3
Ball position relative to the robot’s hip    (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
ballv
3
$\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s
(0,1,2)
(1,-1,1)
balld
1
Euclidean distance from ball to robot’s hip
0
1
targdir
1
Filteredd target direction, relative to the robot’s torso
0
-1
targv
1
$\frac{d\text{targdir}}{dt}\,$ for $\Delta t=0.02$ s
0
-1
a A limit prevents the neural network from encountering higher values during testing than during training. The value 96 was optimized as a hyperparameter, as it influences the final behavior.
b Values extracted from 6D pose [4]
c The symmetry operation associated with this sagittal reflection depends on the joints’ order
d targdir is limited to 80 deg and a maximum variation of 20 deg per time step (0.02 s)
@@
1. Action
Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter
An exponential moving average filter is applied with a smoothing factor of 0.15
3. Global scale
A global scaling parameter allows the robot to gradually slow down and revert to the Step Baseline primitive for an easy transition to walking
4. Map
The individual values are mapped to the desired initial exploration range
5. Primitive
Add Step Baseline primitive
6. Constraints
Limit leg joints to prevent self-collisions
7. IK
Apply inverse kinematics to obtain the target joint positions
8. Conversion
First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.
@@ 8. Conversion
Reset
The robot is initially placed near the ball with both a random position and orientation. The ball is given a random velocity with a magnitude inferior to 1 m/s. Then, the robot aligns itself with the ball and starts the approach until it is close enough ($<$ 0.25 m).
Termination
Robot has fallen ($z<0.40$), timeout (40 s), ball was lost, or ball cannot be seen
Episode
A new target direction is randomly generated on average every 0.33 s
Reward
@@
step
1
Step counter (max: 96)a [3]
0
1
z
1
Head heightb
0
1
dz
1
$\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s
0
1
roll
1
Torso roll angleb
0
-1
pitch
1
Torso pitch angleb
0
1
gyro
3
Gyroscope (x-forward, y-left, z-up)
(0,1,2)
(-1,1,-1)
acc
3
Accelerometer (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
feet
12
(l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
(6,7,8,9,10,11, 0,1,2,3,4,5)
(1,-$,1$,-1,1,1-1,1)
joints
20
All joint positions excluding head and toes
Sagittal reflectionc
Sagittal reflectionc
djoints
20
$\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s
Sagittal reflectionc
Sagittal reflectionc
sbprog
1
Step Baseline cycle progress $\in[0,1]$
0
1
sbleg
2
One-hot encoding for the current Step Baseline swing leg (left leg, right leg)
(1,0)
(1,1)
sbsin
1
Step Baseline oscillation: sin(sbprog ×$\times\pi$)
0
1
ballp
3
Ball position relative to the robot’s hip    (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
ballv
3
$\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s
(0,1,2)
(1,-1,1)
balld
1
Euclidean distance from ball to robot’s hip
0
1
targp
2
2D target position vector, relative to the robot’s torso (x-forward, y-left)
(0,1)
(1,-1)
opp
2
Weightedd average of closest opponent’s visible limbs positions, relative to our robot’s torso
(0,1)
(1,-1)
a A limit prevents the neural network from encountering higher values during testing than during training. The value 96 was optimized as a hyperparameter, as it influences the final behavior.
b Values extracted from 6D pose [4]
c The symmetry operation associated with this sagittal reflection depends on the joints’ order
d The relative weight for each limb is $w_{r}=10^{-6d}$, where $d$ is the respective Euclidean distance
@@
1. Action
Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter
An exponential moving average filter is applied with a smoothing factor of 0.2
3. Map
The individual values are mapped to the desired initial exploration range
4. Primitive
Add Step Baseline primitive
5. Constraints
Limit leg joints to prevent self-collisions
6. IK
Apply inverse kinematics to obtain the target joint positions
7. Conversion
First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.

### v2

@@ 4. Map The raw values are mapped to the position ranges of each individual joint
Reset The robot is reset to a fixed initial pose with bent knees, which is then added as a continuous bias (see Table 11, 5a)
Termination Robot has fallen ($z<0.27$) or timeout (4 s)
Episode There is no specific configuration during the episode
Reward $r=\frac{d\text{x}}{dt}\,$ for $\Delta t=0.06$ s. There is no penalty for deviations in $y$. The reward is updated every vision cycle.
@@
Reset Sprint Forward is executed for 1.4 s (this parameter was found to be the minimal period for which the resulting behavior, Sprint & Rotate, achieves maximum speed)
Termination Robot has fallen ($z<0.34$) or timeout (7 s)
Episode A new target direction is randomly generated every 0.28 s (cycle caption period, see Fig. 3). The target’s range is [-10,10] deg, relative to the current orientation of the robot’s torso.
@@
Reset The ball is placed in a random but reachable position. Sprint Forward is executed for 1.4 s. Then, Sprint & Rotate is executed until the ball is within 1 meter.
Termination Robot has fallen ($z<0.33$) or timeout (0.4 s)
Episode There is no specific configuration during the episode
@@
Table 15: Sprint-Kick hyperparam.: Sprint Forward (SF), Sprint & Rotate (SR), Kick (K)
@@ A.3 Locomotion set
step 1 Step counter (max: 120) [^1] A limit prevents the neural network from encountering higher values during testing than during training. The value 120 was optimized as a hyperparameter, as it influences the final behavior. [3] 0 1
z 1 Head height [^2] Values extracted from 6D pose [5] 0 1
dz 1 $\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s 0 1
roll 1 Torso roll angle [^2] Values extracted from 6D pose [5] 0 -1
pitch 1 Torso pitch angle [^2] Values extracted from 6D pose [5] 0 1
gyro 3 Gyroscope (x-forward, y-left, z-up) (0,1,2) (-1,1,-1)
acc 3 Accelerometer (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
feet 12 (l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz) (6,7,8,9,10,11, 0,1,2,3,4,5) (1,-$,1$,-1,1,1,-1,1)
ankles 6 (l)eft and (r)ight ankle position relative to the robot’s hip: (lx,ly,lz,rx,ry,rz) (3,4,5,0,1,2) (1,-1,1,1,-1,1)
feetori 6 (l)eft and (r)ight feet orientation relative to the robot’s hip: (lx,ly,lz,rx,ry,rz) (3,4,5,0,1,2) (-1,1,-1,-1,1,-1)
shldrs 4 (l)eft and (r)ight shoulder joints positions Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
action 16 Action array generated in last time step Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
sbprog 1 Step Baseline cycle progress $\in[0,1]$ 0 1
sbleg 2 One-hot encoding for the current Step Baseline swing leg (left leg, right leg) (1,0) (1,1)
targp 2 Filtered444targp is limited to a magnitude of 0.5 m and a maximum variation of 0.014 m per time step (0.02 s), and targdir is limited to 45 deg and a maximum variation of 1.6 deg per time step 2D target position vector, relative to the robot’s torso (x-forward, y-left) (0,1) (1,-1)
targv 2 $\frac{d}{dt}\begin{bmatrix}\text{targp}_{x}&\text{targp}_{y}\end{bmatrix}$ for $\Delta t=0.02$ s (0,1) (1,-1)
targdir 1 Filtered444targp is limited to a magnitude of 0.5 m and a maximum variation of 0.014 m per time step (0.02 s), and targdir is limited to 45 deg and a maximum variation of 1.6 deg per time step target direction, relative to the robot’s torso 0 -1
@@
1. Action Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter An exponential moving average filter is applied with a smoothing factor of 0.2
3. Global scale The filtered action is multiplied by $\;\|\text{targp}\|_{2}\times 3.5+0.3\;$ if $\;\|\text{targp}\|_{2}<0.2\;$ to encourage the robot to minimize movement as it approaches the target position
4. Map The individual values are mapped to the desired initial exploration range
5. Primitive Add Step Baseline primitive
6. Constraints Limit leg joints to prevent self-collisions
7. IK Apply inverse kinematics to obtain the target joint positions
8. Conversion First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.
@@ 7. IK Apply inverse kinematics to obtain the target joint positions
Reset Execute Get Up behavior
Termination Robot has fallen ($z<0.35$)
Episode Two targets are created: position (TP) and direction (TD). A random velocity is periodically assigned (avg. interval of 1.6 s) to a moving TP within a work area of 196 $\text{m}^{2}$. Upon leaving the work area, the TP is relocated to its center. In 30% of episodes, TD coincides with TP, whereas in 70% of episodes, TD changes randomly at an avg. interval of 4 s.
@@
Param. name Data size $\times$32b FP Description
step 1 Step counter [^1] Step counter limits are not needed when testing episodes are never longer than training episodes [3]
z 1 Head height [^2] Values extracted from 6D pose [5]
dz 1 $\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s
roll 1 Torso roll angle [^2] Values extracted from 6D pose [5]
pitch 1 Torso pitch angle [^2] Values extracted from 6D pose [5]
gyro 3 Gyroscope (x-forward, y-left, z-up)
acc 3 Accelerometer (x-forward, y-left, z-up)
feet 12 (l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
joints 16 All joint positions excluding head, toes and 4 elbow joints
djoints 16 $\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{15}\end{bmatrix}$ for $\Delta t=0.02$ s
ballp 3 Ball position relative to the robot’s hip (x-forward, y-left, z-up)
ballv 3 $\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s
balld 1 Euclidean distance from ball to robot’s hip
targdir 1 Target direction, relative to the robot’s torso
targd  [^3] The target distance is only included in Short Kick 1 Target distance
@@
1. Action Raw target speed for all joints, excluding head, toes and 4 elbow joints (16 values)
2. Map The individual values are mapped to the desired initial exploration range
3. Bias An initial bias is added to pull the kicking leg back for 0.04 s (for Short Kick) or 0.12 s (for Long Kick) to generate a back swing
@@ 2. Map The individual values are mapped to the desired initial exploration range
Reset The target direction is constant. The robot is initially placed near the ball with both a random position and orientation. The ball is given a random velocity with a magnitude inferior to 1 m/s. Then, the robot aligns itself with the ball and the target as it approaches. The alignment is stricter for the Long Kick. Additionally, for the Short Kick, the target distance is randomly set between 3 m and 9 m.
Termination Timeout (0.20 s for Short Kick and 0.32 s for Long Kick)
Episode There is no specific configuration during the episode
@@
step 1 Step counter (max: 96) [^1] A limit prevents the neural network from encountering higher values during testing than during training. The value 96 was optimized as a hyperparameter, as it influences the final behavior. [3] 0 1
z 1 Head height [^2] Values extracted from 6D pose [5] 0 1
dz 1 $\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s 0 1
roll 1 Torso roll angle [^2] Values extracted from 6D pose [5] 0 -1
pitch 1 Torso pitch angle [^2] Values extracted from 6D pose [5] 0 1
gyro 3 Gyroscope (x-forward, y-left, z-up) (0,1,2) (-1,1,-1)
acc 3 Accelerometer (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
feet 12 (l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz) (6,7,8,9,10,11, 0,1,2,3,4,5) (1,-$,1$,-1,1,1,-1,1)
joints 20 All joint positions excluding head and toes Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
djoints 20 $\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
sbprog 1 Step Baseline cycle progress $\in[0,1]$ 0 1
sbleg 2 One-hot encoding for the current Step Baseline swing leg (left leg, right leg) (1,0) (1,1)
sbsin 1 Step Baseline oscillation: sin(sbprog ×$\times\pi$) 0 1
ballp 3 Ball position relative to the robot’s hip    (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
ballv 3 $\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s (0,1,2) (1,-1,1)
balld 1 Euclidean distance from ball to robot’s hip 0 1
targdir 1 Filtered444targdir is limited to 80 deg and a maximum variation of 20 deg per time step (0.02 s) target direction, relative to the robot’s torso 0 -1
targv 1 $\frac{d\text{targdir}}{dt}\,$ for $\Delta t=0.02$ s 0 -1
@@
1. Action Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter An exponential moving average filter is applied with a smoothing factor of 0.15
3. Global scale A global scaling parameter allows the robot to gradually slow down and revert to the Step Baseline primitive for an easy transition to walking
4. Map The individual values are mapped to the desired initial exploration range
5. Primitive Add Step Baseline primitive
6. Constraints Limit leg joints to prevent self-collisions
7. IK Apply inverse kinematics to obtain the target joint positions
8. Conversion First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.
@@ 7. IK Apply inverse kinematics to obtain the target joint positions
Reset The robot is initially placed near the ball with both a random position and orientation. The ball is given a random velocity with a magnitude inferior to 1 m/s. Then, the robot aligns itself with the ball and starts the approach until it is close enough ($<$ 0.25 m).
Termination Robot has fallen ($z<0.40$), timeout (40 s), ball was lost, or ball cannot be seen
Episode A new target direction is randomly generated on average every 0.33 s
@@
step 1 Step counter (max: 96) [^1] A limit prevents the neural network from encountering higher values during testing than during training. The value 96 was optimized as a hyperparameter, as it influences the final behavior. [3] 0 1
z 1 Head height [^2] Values extracted from 6D pose [5] 0 1
dz 1 $\frac{d\text{z}}{dt}\,$ for $\Delta t=0.04$ s 0 1
roll 1 Torso roll angle [^2] Values extracted from 6D pose [5] 0 -1
pitch 1 Torso pitch angle [^2] Values extracted from 6D pose [5] 0 1
gyro 3 Gyroscope (x-forward, y-left, z-up) (0,1,2) (-1,1,-1)
acc 3 Accelerometer (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
feet 12 (l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz) (6,7,8,9,10,11, 0,1,2,3,4,5) (1,-$,1$,-1,1,1,-1,1)
joints 20 All joint positions excluding head and toes Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
djoints 20 $\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^3] The symmetry operation associated with this sagittal reflection depends on the joints’ order
sbprog 1 Step Baseline cycle progress $\in[0,1]$ 0 1
sbleg 2 One-hot encoding for the current Step Baseline swing leg (left leg, right leg) (1,0) (1,1)
sbsin 1 Step Baseline oscillation: sin(sbprog ×$\times\pi$) 0 1
ballp 3 Ball position relative to the robot’s hip    (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
ballv 3 $\frac{d}{dt}\begin{bmatrix}\text{ball}_{x}&\text{ball}_{y}&\text{ball}_{z}\end{bmatrix}$ for $\Delta t=0.04$ s (0,1,2) (1,-1,1)
balld 1 Euclidean distance from ball to robot’s hip 0 1
targp 2 2D target position vector, relative to the robot’s torso (x-forward, y-left) (0,1) (1,-1)
opp 2 Weighted [^4] The relative weight for each limb is $w_{r}=10^{-6d}$, where $d$ is the respective Euclidean distance average of closest opponent’s visible limbs positions, relative to our robot’s torso (0,1) (1,-1)
@@
1. Action Relative 6D pose of ankles relative to hip (12 values), shoulder joint positions (4 values)
2. IIR Filter An exponential moving average filter is applied with a smoothing factor of 0.2
3. Map The individual values are mapped to the desired initial exploration range
4. Primitive Add Step Baseline primitive
5. Constraints Limit leg joints to prevent self-collisions
6. IK Apply inverse kinematics to obtain the target joint positions
7. Conversion First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.